# medDreamer: Model-Based Reinforcement Learning with Latent Imagination on Complex EHRs for Clinical Decision Support

### v2

@@ 4.2.3 RQ3: Is the learned value function predictive of patient survival?
Model WPDIS CWPDIS
No AFI 0.20 0.14
MDP1 Only 0.19 0.14
MDP2 Only 0.27 0.19
E2E 0.21 0.13
ActorCritic 0.23 0.17
Curriculum 0.23 0.17
@@
[…]
@@ Appendix B Features and Actions
Sepsis observations contain 40 physiological measurements, including demographics, vital signs, and laboratory test results, listed in Table 3. For the action space, two interventions are included: IV fluids and Vasopressors. We discretize the dosage of each treatment into four quantile bins along with a no-treatment option, resulting in five discrete levels per treatment. Consequently, the joint action space consists of 5×5=255$\times$ 5=255 × 5 = 25 possible treatment combinations.
@@
PEEP (cmH2O) Low 5$\leq$ 5≤ 5
High >5>5> 5
FiO2 (%) Low <35<35< 35
Medium 353535505050
High 50$\geq$ 50≥ 50
Tidal Volume (ml/kg IBW) Low <6.5<6.5< 6.5
Medium 6.56.56.5888
High 8$\geq$ 8≥ 8
@@ Appendix D Mortality vs Return
Appendix E Broader Impact

### v3

@@ 4.2.3 RQ3: Is the learned value function predictive of patient survival?
Model Sepsis MV
@@
No AFI 9.10 0.20 0.14 15.77 0.29 0.18
MDP1 Only 7.78 0.19 0.14 17.11 0.29 0.22
MDP2 Only 6.81 0.21 0.13 14.27 0.31 0.22
E2E 10.11 0.21 0.13 8.22 0.26 0.16
ActorCritic 7.65 0.23 0.17 19.81 0.31 0.23
Curriculum 7.65 0.23 0.17 15.93 0.29 0.21
medDreamer($\tau=5$) 8.34 0.26 0.19 21.99 0.29 0.17
MedDreamer($\tau=10$) 9.24 0.27 0.19 24.06 0.32 0.25
medDreamer($\tau=30$) 8.81 0.21 0.15 8.21 0.29 0.18
medDreamer($\tau=50$) 8.42 0.24 0.17 18.21 0.28 0.17
medDreamer($H=5$) 8.66 0.24 0.17 16.91 0.29 0.2
MedDreamer($H=15$) 9.24 0.27 0.19 24.06 0.32 0.25
medDreamer($H=25$) 9.08 0.26 0.19 16.92 0.28 0.21
medDreamer($H=35$) 9.01 0.24 0.17 16.15 0.28 0.2
LSTM 8.89 0.22 0.16 16.58 0.29 0.19
Transformer 9.11 0.27 0.19 16.90 0.29 0.20
Random mask 9.17 0.25 0.19 18.16 0.26 0.16
@@ Appendix B Features and Actions
Sepsis observations contain 40 physiological measurements, including demographics, vital signs, and laboratory test results, listed in Table 3. For the action space, two interventions are included: IV fluids and Vasopressors. We discretize the dosage of each treatment into four quantile bins along with a no-treatment option, resulting in five discrete levels per treatment. Consequently, the joint action space consists of $5\times 5=25$ possible treatment combinations.
@@
PEEP (cmH2O) Low $\leq 5$
High >5>5
FiO2 (%) Low <35<35
Medium 35355050
High $\geq 50$
Tidal Volume (ml/kg IBW) Low <6.5<6.5
Medium 6.56.588
High $\geq 8$
@@ Appendix D Mortality vs Return
Appendix E Additional results
E.1 Model Robustness Analysis
The missingness observed in the dataset is intrinsic, largely representing informative missingness where the absence of a measurement conveys clinical significance. As the dataset lacks explicit labels to differentiate informative from random missingness, we evaluated model robustness by simulating increased data sparsity.We applied random masks to an additional 10%, 20%, and 50% of the observed values. As demonstrated in Table 6, the model retains strong performance on the Sepsis task with up to 20% induced missingness.
Table 6: Robustness experiment results on Sepsis task with varying mask percentages.
Mask% WIS WPDIS CWPDIS
0 9.24 0.27 0.19
10 9.23 0.25 0.19
20 9.17 0.25 0.19
50 8.44 0.22 0.15
E.2 Evaluation on Renal Replacement Therapy
To assess the generalizability of medDreamer across distinct patient populations and healthcare systems, we extended our evaluation to the AmsterdamUMCdb, a European ICU dataset. This analysis focused on the task of renal replacement therapy, where the agent is required to determine the necessity of dialysis and select the appropriate mode. It is important to note that the OPE metrics for this task are negative, reflecting a reward function design that incorporates a higher density of negative values compared to previous tasks. As presented in Table 7, medDreamer consistently outperforms both the clinician and the baseline Dreamer model, validating the framework’s adaptability to non-US datasets and distinct clinical decision-making tasks.
Table 7: Performance comparison on AmsterdamUMCdb for the Renal Replacement Therapy task.
Method WIS WPDIS CWPDIS
Clinician -0.67 -0.08 -0.03
dreamer -0.82 -0.07 -0.03
medDreamer -0.18 -0.06 -0.02
E.3 Comparison with Additional Baselines
We incorporated two additional baseline model results on the Sepsis dataset: DAG and MeDT. The comparative results are shown in Table 8. Compared to DAG and MeDT, our method’s designed AFI module endows our model with better performance on irregular temporal data. At the same time, our two-phase training strategy ensures safety under the clinical context.
Table 8: Comparison with additional baselines on the Sepsis dataset.
Method WIS WPDIS CWPDIS
Random 2.17 -0.01 -0.01
Clinician 4.87 0.14 0.09
EZ-Vent[liu2024reinforcement](BCQ) 10.13 0.02 0.07
DeepVent[kondrup2023towards](CQL) 10.42 0.03 0.07
DDQN[raghu2017deep] 10.10 0.02 0.07
DAG[yin2022deconfounding] 6.12 0.03 0.01
MeDT[rahman2024empowering] 7.63 0.13 0.06
MBRL-BNN[raghu2018model] 4.90 0.00 0.01
MBRL-GMM[li2022electronic] -0.74 -0.01 -0.01
Dreamer[hafner2024masteringdiversedomainsworld] 6.85 0.20 0.13
medDreamer 9.25 0.27 0.19
Appendix F Broader Impact